# New theoretical Fe II templates for bright quasars

### v1

@@ Abstract
We present a set of new theoretical Fe II templates for bright quasars covering a wavelength range of 1000-10000 Å, based on the recent atomic database available in the C23.00 version of the photoionization code CLOUDY. We compute a grid of models for a range of incident photon flux and gas density and for multiple microturbulence velocities. We analyze the ratios of Fe II emission over a variety of wavebands and compare them with observations. Our key results are: (1) Despite the use of the newest atomic data we still confirm the long-standing problem that the predicted Fe II UV/optical ratio is significantly larger than that observed in the AGN spectra. (2) The ratio is not significantly affected by the variations in the microturbulence and the metallicity. (3) The turbulence can create an additional apparent velocity shift of up to 1000 km/s in the spectra. (4) There is no single Fe II template that can fit the observational data covering UV to optical wavelength range. We shortly discuss the most likely effects responsible for the Fe II UV/optical mismatch problem: the assumption of the constant column density and the assumption of the isotropic emission implying equal contribution of the bright irradiated faces and the dark shielded faces of the clouds.
@@ 1 Introduction
The dominant component of the bright active galactic nuclei (AGN) in the optical/UV band is the continuum component coming from the cold optically thick geometrically thin disk (see e.g. Krolik, 1999; Karas et al., 2021; Zajaček et al., 2023a, for reviews). Locally this component is well described as a power law although at high quality and/or in broadband data some curvature of the component is visible (e.g. Capellupo et al., 2015). Superimposed on this continuum component are numerous strong and broad emission lines like H, Mg II, C IV, and Ly , depending on the redshift range covered by the data. These lines are easy to identify in the individual spectra of AGN and quasar composite spectra (e.g. Francis et al., 1991; Vanden Berk et al., 2001; Telfer et al., 2002; Scott et al., 2004; Ren et al., 2022). These lines come from the Broad Line Region (BLR) which is located at a fraction of a parsec from the central supermassive black hole (SMBH). This region reprocesses the bulk of the radiation from the central parts very close to the SMBH.
Apart from these well-resolved strong individual lines, we have a contribution from transitions in weakly ionized heavy elements, predominantly iron. These transitions are so numerous that they form a pseudo-continuum well visible in the near-IR (NIR), optical and UV bands. This pseudo-continuum affects considerably the fitting of the strong individual emission lines from BLR. The issue has been identified a long time ago (Sargent, 1968; Collin-Souffrin et al., 1979; Wills et al., 1985, 1980). The important role of Fe II diagnostics has been recognized by Boroson & Green (1992) in their studies of the many AGN parameters using the Principal Component Analysis (PCA) approach. Their study led to the identification of the key role of the Eigenvector 1 (EV1) in the quasar sequence. The ratio, , of the equivalent width (EW) of the Fe II in the 4434-4684 Å wavelength range to the EW of H played a key role in EV1, and later it became the foundation of the optical Quasar Main Sequence plots showing the Full Width Half Maximum (FWHM) of H vs (Sulentic et al., 2000; Shen & Ho, 2014; Marziani et al., 2018; Panda et al., 2018, 2019b).
Since Fe II transitions are so numerous, certain standard templates were developed separately for the optical and for the UV bands which were used for detailed modeling of AGN spectra. Some of the templates were purely observational, i.e. created by selecting the object with relatively narrow lines and subtracting the known lines and continuum. The residual was considered as coming from Fe II (e.g. Vestergaard & Wilkes, 2001; Marziani et al., 2003; Park et al., 2022). Other templates were mostly theoretical, based on computing the transition rates (e.g. Bruhweiler & Verner, 2008; Kovačević-Dojčinović & Popović, 2015). Some templates were a combination of both approaches (Tsuzuki et al. 2006, but also partially Kovačević-Dojčinović & Popović 2015).
@@
In this paper, we use the latest version C23.00 of the photoionization code CLOUDY (Chatzikos et al., 2023), which contains a large database of the Fe II transitions, to develop a new theoretical Fe II template that covers a wide wavelength range of 1000-10000 Å. This new template will be useful to fit the AGN spectra simultaneously from optical to UV bands. Some properties of Fe II emission, like the ratio of the Fe II to Mg II line in UV and the trends with the EV1 were studied by Sarkar et al. (2021); Dias dos Santos et al. (2024) but here we study systematically several correlations which can also be tested against observational data in the NIR to UV wavebands.
@@ 2 Method
We computed several grids of photoionization models that span five decades in the hydrogen gas density, 9 14, and the incident hydrogen-ionizing photon flux, 17 22, with a step size of 0.25 dex. We adopted a fixed hydrogen column density of 10 cm and set the abundance to the default solar abundance, as in Sarkar et al. (2021). For the description of the spectral shape of the incident radiation, we used the standard AGN.SED available in the CLOUDY database, which is similar to the continuum of Mathews & Ferland (1987).
@@
We generated such models with different values of microturbulence velocity (V). We used six values of microturbulence velocity: 0, 10, 20, 30, 50, and 100 km/s. Such range for the microturbulence velocity was inspired by previous works (Baldwin et al., 2004; Bruhweiler & Verner, 2008; Panda et al., 2018). For a certain set of values of and , we did not get any Fe II emission. We listed these values in Table 1.
Having the set of computed spectra, we constructed several ratios of Fe II emission by integrating the spectra over the selected wavelength ranges. Such ratios are frequently studied in the observed spectra, and we aim to analyze the trends in these ratios as functions of the density and the ionizing flux.
We also examine the effect of metallicity on these ratios by varying metallicity as 2, 5, 7, 10 and 20 times the solar metallicity.
21.25 | 9.00 |
21.50 | 9.00 |
21.50 | 9.25 |
21.75 | 9.00 |
21.75 | 9.25 |
21.75 | 9.50 |
22.00 | 9.00 |
22.00 | 9.25 |
22.00 | 9.50 |
22.00 | 9.75 |
Note: This set of parameters is valid for all the turbulence velocities except for zero. In case of no turbulence velocity in addition to the upper parameters, Fe II emission was also not obtained for = 22 and = 10.
2.1 Definitions of optical and UV Fe II
Since the Fe II emission is of a broad band character the integrated value depends critically on the adopted range of integration. In our computations, we use several such ranges that were frequently considered in the literature.
@@ 2.1 Definitions of optical and UV Fe II
To characterize the broadband emission in the optical range we integrate Fe II from 3500 to 7200 Å (Véron-Cetty et al., 2004). Some of the studies concentrate on much narrower ranges. For example, the range from 4434 - 4684 Å was introduced to parameterize the Fe II emission close and bluewards of the H line and to characterize the relative importance of the Fe II to H emission by the corresponding ratio (Boroson & Green, 1992). This ratio was well studied observationally in many papers (e.g. Osterbrock, 1977; Bergeron & Kunth, 1984; Shen & Ho, 2014; Śniegowska et al., 2018; Martínez-Aldama et al., 2021; Marziani et al., 2022).
@@
Broad UV Fe II range can be well characterized by the Fe II emission integrated in the wavelength range from 1250 to 3090 Å (Vestergaard & Wilkes, 2001). This range contains the Fe II bump at 2500 Å seen in the spectra as well as in early models of Fe II emission (Wills et al., 1985).
@@
We also conduct a systematic analysis of two UV and two optical wavelength ranges that may be used to determine the spectral shape. The UV ranges, 2650-2800 Å and 2800-3090 Å characterize the Fe II emission on each side of the Mg II line. The optical ranges are 4434-4684 Å and 5150-5350 Å corresponding to the Fe II emission at the blue and red sides of the H line, respectively. The Fe II emission on both sides of Mg II and H forms clear local bumps or peaks. Therefore, for convenience, we refer to these regions as blue and red wings (e.g. Sarkar et al., 2021).
@@ 3 Results
3.1 UV to optical Fe II emission
We calculated the ratio of the UV to optical broadband Fe II emission with the integration range specified in Section 2.1. The results are shown in Figure 1 using contour plots. The representative value in the central parts of the plot is 0.7 in the log space, i.e. the ratio is typically of the order of 5 in linear space. As already argued by Joly (1981) and Wills et al. (1985), a ratio of the order of 1 in the relative intensities of the optical to UV lines implies a column density of about to cm to balance Fe II line optical depths and the Balmer opacity. This value is only weakly sensitive to the ionization flux or density. However, we see that this ratio rises rapidly when either the ionization parameter becomes very high or very low, and the cloud density rises. At higher values of the ionization parameter, above cm s, this steep rising wall forms a linear trend in the diagram, with the characteristic position correlated with the density. The ratio at this wall exceeds 100. At low values of the ionization flux, below cm s, the ratio depends mostly on the density, changing from the factor 2.5 in the linear scale for densities below cm to values above 100 for densities above cm. Lines are almost vertical there, so for low values of the ionization flux this ratio can serve as a density indicator. At intermediate values of the ionization flux, above cm s, the dependence on the two parameters is more complex. Thus, when we do not have an estimate of the ionization parameter in the source, high values of the UV to optical Fe II emission can indicate either very high or very low ionization flux. Sarkar et al. (2021) stressed the role of the turbulent velocity in the Fe II intensity but the global ratio of the UV to optical is relatively unaffected.
The ratio shown in Figure 1 represents the expectations based on the assumption of a single zone responsible for both UV and optical emission, which may not be the case in actual objects. We address this issue in the discussion.

### v2

@@ Abstract
We present a set of new theoretical Fe II templates for bright quasars covering a wavelength range of 1000-10000 Å based on the recent atomic database available in the C23.00 version of the photoionization code CLOUDY. We compute a grid of models for a range of incident photon flux, gas density, and multiple microturbulence velocities. We examine the equivalent widths (EWs) and the ratios of Fe II emission over various wavebands and compare them with observations. Our key results are: (1) The flux generated from the shielded side of the cloud is insufficient to describe the measured Fe II emission. (2) Despite using the newest atomic data we still confirm the long-standing problem that the predicted Fe II UV/optical ratio is significantly larger than that observed in the AGN spectra. (3) The Fe II UV/optical ratio is not significantly affected by the variations in the microturbulence and the metallicity. (4) The microturbulence can create an additional apparent velocity shift of up to 1000 km/s in the spectra. (5) There is no Fe II template based on a single set of physical parameters that can fit the observed UV to optical Fe II emission spectra. We shortly discuss the most likely effects responsible for the Fe II UV/optical mismatch problem: the assumption of the constant density clouds and the heating mechanism for Fe II emitting clouds.
@@ 1 Introduction
The dominant component of the bright active galactic nuclei (AGN) in the optical/UV band is the continuum component coming from the cold optically thick geometrically thin disk (see e.g. Krolik, 1999; Karas et al., 2021; Panda et al., 2023; Zajaček et al., 2024a, 2025, for reviews). Locally this component is well described as a power law although at high quality and/or in broadband data some curvature of the component is visible (e.g. Capellupo et al., 2015). Superimposed on this continuum component are numerous strong and broad emission lines like H, Mg II, C IV, and Ly , depending on the redshift range covered by the data. These lines are easy to identify in the individual spectra of AGN and quasar composite spectra (e.g. Francis et al., 1991; Vanden Berk et al., 2001; Telfer et al., 2002; Scott et al., 2004; Ren et al., 2022). These lines come from the Broad Line Region (BLR) which is located at a fraction of a parsec from the central supermassive black hole (SMBH). This region reprocesses the bulk of the radiation from the central parts very close to the SMBH.
Apart from these well-resolved strong individual lines, we have a contribution from transitions in weakly ionized heavy elements, predominantly iron. These transitions are so numerous that they form a pseudo-continuum well visible in the near-IR (NIR), optical and UV bands. This pseudo-continuum affects considerably the fitting of the strong individual emission lines from BLR. The issue has been identified a long time ago (Sargent, 1968; Collin-Souffrin et al., 1979; Wills et al., 1985, 1980). The important role of Fe II diagnostics has been recognized by Boroson & Green (1992) in their studies of the many AGN parameters using the Principal Component Analysis (PCA) approach. Their study led to the identification of the key role of the Eigenvector 1 (EV1) in the quasar sequence. The ratio, , of the equivalent width (EW) of the Fe II in the 4434-4684 Å wavelength range to the EW of H played a key role in EV1, and later it became the foundation of the optical Quasar Main Sequence plots showing the Full Width at Half Maximum (FWHM) of H vs (Sulentic et al., 2000; Shen & Ho, 2014; Marziani et al., 2018; Panda et al., 2018, 2019b; Panda & Śniegowska, 2024).
Since Fe II transitions are so numerous, certain standard templates were developed separately for the optical and for the UV bands which were used for detailed modeling of AGN spectra. Some of the templates were purely observational, i.e. created by selecting the object with relatively narrow lines and subtracting the known lines and continuum. The residual was considered as coming from Fe II (e.g. Vestergaard & Wilkes, 2001; Marziani et al., 2003; Park et al., 2022; Marziani et al., 2021a). Other templates were mostly theoretical, based on computing the transition rates (e.g. Bruhweiler & Verner, 2008; Kovačević-Dojčinović & Popović, 2015; Panda, 2021a). Some templates were a combination of both approaches (Tsuzuki et al. 2006, but also partially Kovačević-Dojčinović & Popović 2015).
@@
Recently, Sarkar et al. (2021) incorporated three extensive Fe II atomic data sets to the development version of the photoionization code CLOUDY (Ferland et al., 2017) and examined the predicted Fe II spectra in the UV and optical regions for different SED shapes. They compared their model-predicted Fe II spectra with the observed UV (Vestergaard & Wilkes, 2001) and optical (Véron-Cetty et al., 2004) Fe II templates of I Zw 1 Seyfert galaxy to constrain the properties of Fe II emitting gas in the BLR. They concluded that the model predicted Fe II spectra for a BLR cloud having gas density cm-3, ionizing flux cm-2 s-1, and microturbulence V km s-1 was consistent with the observed Fe II templates. Additionally, they investigated the (Fe II)/(Mg II) line intensity ratio and found that the ratio predicted by their best-fitted model with solar abundance matched the observed value.
In this paper, we use the latest version C23.00 of CLOUDY (Chatzikos et al., 2023), which contains all new Fe II atomic databases, to develop a set of new theoretical Fe II templates that cover a wide wavelength range of 1000-10000 Å. These new templates will be useful to fit the AGN spectra simultaneously from optical to UV bands. We analyze the strength of Fe II emission across various commonly used wavebands in both the UV and optical parts of the spectrum, as well as the ratios of Fe II fluxes between these wavebands. Additionally, we assess the equivalent widths of Fe II blends for both the illuminated and shielded parts of the BLR cloud. To understand the role of microturbulence, we explore how varying microturbulence velocities influence the Fe II spectra, focusing on their effect on different Fe II flux ratios. For a fixed value of microturbulence, we further investigate the impact of changing metallicity on the Fe II spectra. Finally, we try to constrain the physical parameters of the BLR gas producing Fe II emission by comparing our model-predicted Fe II templates with the observed spectra of quasars RM 102 and 1H 0413-4031. While Sarkar et al. (2021) tested their model-predicted Fe II spectra exclusively on the narrow-line Seyfert 1 galaxy I Zw 1, our objects are more representative of typical quasars.
@@ 2 Method
We computed several grids of photoionization models that span five decades in the hydrogen gas density, 9 14, and the incident hydrogen-ionizing photon flux, 17 22, with a step size of 0.25 dex. We adopted a fixed hydrogen column density of 1024 cm-2 and set the abundance to the default solar abundance, as in Sarkar et al. (2021). We use the assumption of a constant total hydrogen gas density of the clouds for our photoionization computations.
The shape of the incident continuum SED plays a crucial role in predicting the Fe II spectra in AGN (e.g. Sarkar et al., 2021). The equivalent widths (EWs) of the Fe II UV and optical multiplets, as well as their ratios, are strongly affected by the shape of the incident SED (see Figure 11 of Sarkar et al. (2021). Since we plan to compare our model-predicted Fe II spectra with the observational data of quasar RM 102, which has an Eddington ratio 0.33 (see Section 2.1), we adopted the intermediate SED of Jin et al. (2012) which is suitable for sources with for our CLOUDY simulations.
@@
We generated such models with different values of microturbulence velocity. We used four values of microturbulence velocity: 0, 20, 50, and 100 km/s. Such a range for the microturbulence velocity was inspired by previous works (Baldwin et al., 2004; Bruhweiler & Verner, 2008; Panda et al., 2018; Panda, 2021b). We also examine the effect of metallicity on the Fe II emission by varying metallicity as 2, 5, and 10 times the solar metallicity.
2.1 Observational data
We provide very preliminary tests of the new templates with the observational data. As for the basic object for these tests, we selected the quasar SDSS J141352.99+523444.2 (SDSS RM102) with RA=, Dec=, =0.861140.00029, log(L3000[erg/s])= 45.00.0005, log(L5100[erg/s])= 45.720.0005, log ( [M⊙])=8.0860.036, and an Eddington ratio () of 0.3290.027 (Shen et al., 2019). The spectrum of the source covers a broad wavelength range (2000-5500 Å). This allowed us to cover both the rest-frame optical Fe II as well as UV Fe II and to test the feasibility of using the same Fe II template to fit the entire UV-to-optical spectral range.
The size of its BLR region has been estimated by different methods, in this paper we took the one estimated by the Interpolated Cross-Correlation Function (ICCF), and days in the observed-frame (Shen et al., 2024). Using the SDSS Science Archive Server (SAS) from the DR16111https://dr16.sdss.org/optical/spectrum/search, we found 70 spectroscopic epochs observed during four years. All the spectra were combined into a median spectrum to obtain a high signal-to-noise spectrum, S/N=30 and 35 at 3000Å and 4000Å, respectively. Due to the presence of spikes around 5100Å, it was not possible to estimate the S/N at this wavelength. The median spectrum was corrected for the Galactic reddening using the extinction laws of Cardelli et al. (1989) and assuming an extinction value of mag in the band.
2.1.1 Spectral decomposition of RM 102
We started with the commonly used approach in which the continuum is modelled as a single power law. Since we could not identify any stellar absorption feature (CaII H and K, G band, Mg I), we simply assumed that the host galaxy contribution is negligible. We considered all the emission lines reported by Vanden Berk et al. (2001) between 2000 and 5500 Å, 31 emission lines in total. We included a model of the Balmer Continuum (BC) plus the High-Order Balmer lines (Bernal et al. in prep.). In the case of the Balmer continuum model, we tested different optical depths ( 1) and electron temperatures (=11000, 15000 K) and implemented the Balmer continuum model proposed by Grandi (1982). Regardless of the optical depth and electron temperature considered, the Balmer Continuum contribution was negligible in the spectral fitting and the residuals were consistent in all the cases. Therefore, we decided to use and K, which are the typical values used in large dataset analysis (e.g. Kovačević et al., 2014). To decrease the number of free parameters, we fixed the same FWHM for each family of emission lines: Balmer lines, Helium transitions, forbidden lines (narrow emission lines), and coronal lines. Flux intensities, FWHM, and shift sum a total of 63 free parameters. We applied Gaussian smearing of FWHM = 4000 km s-1 to the UV and optical Fe II template and a single flux intensity value was taken for the UV and optical wavelengths. In the first attempt, we noticed that it was not possible to fit the UV and optical Fe II contribution with the same flux intensity. Thus, following the prescription by Boroson & Green (1992), we performed a careful fitting of all the other spectral features and what remained can be considered as the Fe II component. Due to the overlapping of the emission lines and the Fe II pseudocontinuum in some ranges, the flux intensity of the emission line would be underestimated. So, we considered the semiempirical UV/optical Fe II templates implemented in the fitting code PyQSOFit (Guo et al., 2018) to constrain the intensity of the emission lines. As we previously noticed, even with these empirical Fe II templates, a different flux intensity value is needed to fit the Fe II contribution. The flux intensity factor has a difference of 3.5 between the UV and the optical Fe II contribution. The final spectral fitting is shown in Figure 22 in the appendix. The potential Fe II contribution is shown at the bottom panel of Figure 22. As a reference, we also plotted the UV (dotted line) and optical (dashed-line) Fe II templates implemented in PyQSOFit. The optical Fe II shows good agreement, while a slight increase is observed in the UV Fe II within the wavelength ranges of 3000-3150 Å and 3500-3650 Å. This increase in the UV Fe II could be attributed to the minimal contribution from the BC in this approach.
We, therefore, performed an alternative analysis by modeling the continuum emission using the standard Shakura & Sunyaev (1973) disk model instead of a power-law. We adopted the black hole mass as given in Section 2.1 and treated the Eddington rate as a free parameter during the fitting. The disk emission exhibits a slightly steeper slope compared to the power law used in the previous approach. To achieve a better fit for the red part of the spectrum, we had to account for some level of contamination from the starlight. For the starlight, we used the synthesis stellar model from Vazdekis et al. (2016) with a solar metallicity and an age of 7.0795 Gyr appropriate for the host galaxy. The normalization of this component was a free parameter. To describe the line emission component, we refitted the line spectrum for a new preliminary shape of the continuum. Later, when fitting the disk and the starlight we did not iterate the line emission, but we allowed for a single parameter which could rescale all the line luminosities while preserving the line profiles and their ratios.
In this approach, we modeled the Balmer continuum following the method of Kovačević et al. (2014). Their method generalizes the model of Grandi (1982), and includes 400 higher-order Balmer lines which considerably contribute to emissions above 3646 Å. We assumed the temperature of the clouds as K, and the optical depth of the Balmer edge equal to 1. The normalization was a free component of the model.
The new improved spectral fitting of RM 102 obtained from this method is shown in Figure 1, where the expected Fe II contribution from RM 102 closely matches the empirical Fe II templates.
To locate the source in the plane, we introduce several specific spectral windows to measure Fe II emission in all of them separately. The motivation for the band choice is later discussed in Section 2.2. These bands are listed in Table 1 and are shown in Figure 1. For comparison, the observed EWs of Fe II in these bands from both spectral fitting methods are provided in Table 1. The EW values obtained using the disk continuum method are slightly lower than those from the power-law method, primarily due to the higher contribution from the BC and the slightly different shape of the continuum. The integrated luminosity of the BC from the second approach is higher than that from the previous decomposition, with the ratio of the BC to H flux equal to 6.15. We also measured three ratios: red-to-blue wings of UV Fe II, red-to-blue blends of optical Fe II, and UV to optical Fe II. The measured ratios from both spectral fitting are given in Table 2. In our further analysis, we used the values of EWs and ratios from our second (disk continuum) spectral fitting method that provides somewhat better fit to the spectra of RM 102. A comparison of the expected Fe II contribution from RM 102 using both spectral fitting method is presented in Figure 23 in the appendix.
2.1.2 HE 0413-4031
. Fe II blend Wavelength range (Å) Observed EW (Å) Observed EW (Å) Power-law fit Disk fit Broadband UV Fe II 2000-3090 290.07 266.56 UV-blue wing 2650-2800 41.09 33.37 UV-red wing 2800-3090 81.22 67.50 Broadband optical Fe II 4000-5350 428.82 348.36 Optical-blue blend 4434-4684 109.50 96.22 Optical-red blend 5150-5350 79.27 65.04
As a complementary test object, we use the quasar HE 0413-4031. This quasar has been observed with the Southern African Large Telescope (SALT) for 13 years to perform the reverberation mapping of the Mg II line and the UV Fe II (Zajaček et al., 2020; Prince et al., 2023; Zajaček et al., 2024b). In this case, only a very narrow range is covered, between 2700 Å and 2900 Å in the quasar rest frame, and the lines are broader than for RM 102.
2.2 Definitions of optical and UV Fe II blends
Since the Fe II emission is of a broadband character the integrated value depends critically on the adopted range of integration. We mostly follow the usually adopted values but with some modifications due to the spectral range available for our test object RM 102.
@@ 2.2 Definitions of optical and UV Fe II blends
To characterize the broadband emission in the optical range we integrate Fe II from 4000 to 5350 Å (Véron-Cetty et al., 2004). Some of the studies concentrate on much narrower ranges. For example, the range from 4434 - 4684 Å was introduced to parameterize the Fe II emission close and bluewards of the H line and to characterize the relative importance of the Fe II to H emission by the corresponding ratio (Boroson & Green, 1992). This ratio was well studied observationally in many papers (e.g. Osterbrock, 1977; Bergeron & Kunth, 1984; Shen & Ho, 2014; Śniegowska et al., 2018; Martínez-Aldama et al., 2021; Marziani et al., 2022; Panda & Śniegowska, 2024; Panda et al., 2024).
@@
Broad UV Fe II range can be well characterized by the Fe II emission integrated in the wavelength range from 1250 to 3090 Å (Vestergaard & Wilkes, 2001). We modified this range shortening it on the UV side since the test object did not cover the spectra below 2000 Å. This range contains the Fe II bump at 2500 Å seen in the spectra as well as in early models of Fe II emission (Wills et al., 1985).
@@
We also conduct a systematic analysis of two UV and two optical wavelength ranges that may be used to determine the spectral shape. The UV ranges, 2650-2800 Å and 2800-3090 Å characterize the Fe II emission on each side of the Mg II line. The optical ranges are 4434-4684 Å and 5150-5350 Å corresponding to the Fe II emission at the blue and red sides of the H line, respectively. The Fe II emission on both sides of Mg II forms clear local bumps or peaks. Therefore, for convenience, we refer to these regions as UV Fe II blue and red wings (e.g. Sarkar et al., 2021). The Fe II emission on the blue and red sides of H forms larger blends, which we call optical blue and optical red blends, respectively. A list of Fe II blends considered in this work is given in Table 1.
Ratio | Numerator range (Å) | Denominator range (Å) | Observed value | Observed value |
Power-law fit | Disk fit | |||
Spike-to-gap | (2280–2430) and (2560-2660) | 2430-2560 | 1.65 | 1.58 |
UV-to-optical | 2000-3090 | 4000-5350 | 2.04 | 2.54 |
UV red-to-blue wings | 2800-3090 | 2650-2800 | 1.98 | 2.02 |
Optical red-to-blue blends | 5150-5350 | 4434-4684 | 0.72 | 0.68 |
@@ 3 Results
We begin by examining the strength of Fe II emission, which is affected by the covering factor representing the fraction of the central flux intercepted by the BLR. This limits the parameter space as some parameter combinations do not result in the effective generation of Fe II even for a 100% covering factor. Next, we discuss the issue of the difference between the observed and model-predicted spectral shape of the UV Fe II emission. We then discuss the different Fe II ratios as described in Section 2.2. These ratios are independent of the issue of the covering factor. This is followed by the discussion of the ratio of the Fe II to H which is broadly used to parameterize the quasar population. Finally, we test the new templates in more detail against our two test objects, showing the exemplary spectral fits.
3.1 Equivalent widths of Fe II blends
For our model-predicted spectra, we estimate the equivalent widths (EWs) of all the Fe II blends, given in Table 1, by first computing the integrated Fe II flux in the corresponding wavelength range and then dividing this integrated flux by the incident continuum at a reference wavelength. To minimize the systematic effects associated with the assumed shape of the incident continuum, we choose 3000 Å and 5100 Å as the reference wavelengths in the UV and optical regime, respectively. Using a similar process, we also compute the EWs of all the blends for the observed spectra of RM 102 and mention them in Table 1. Since Ferland et al. (2009) suggested that the observed Fe II emission could originate from the shielded part of the cloud, in addition to the EW contours of total Fe II emission, we plot EW contours for the inward and outward Fe II emissions in Figures 2-5 for microturbulence velocities 0 and 100 km/s. The EW contours for the microturbulence velocities of 20 and 50 km/s are shown in Appendix B. In our model calculations, we assumed that each model cloud within the density and ionizing photon flux grid has a covering factor of 100% and thereby covers 4 steradians of the sky as seen by the (essentially point) source of ionizing photons. But the covering factor for the BLR cloud has a much smaller value—roughly 10-30% (Baldwin et al., 1995; Korista & Goad, 2019; Panda, 2021b). Thus, the observed EW corresponds to the lower limit of the predicted EW. Therefore, the region consisting of EW higher than the observed EW should only be consistent with the observations.
The colored contours in Figures 2-5 mark the regions in the nH- plane whose clouds predict, assuming a 100% covering factor, at least the measured EWs of the various Fe II blends. The clouds corresponding to the models in the empty regions predict too little Fe II emission line strength in the specific Fe II blend, even for a 100% covering factor. Moreover, since the BLR emission does not originate from a single set of cloud parameters with a single source covering fraction, the EW contours in Figures 2-5 are indicative of the cloud parameters that can significantly contribute to the observed Fe II emission line spectrum. The strength of Fe II emission is enhanced when the microturbulence is high, as seen from the comparison of Figure 2 and Figure 3. However, even with microturbulence 100 km s-1 the outward emission is very weak in the UV. For the broadband and the blue UV wing even with the assumption of a 50% covering factor no allowed region appears for outward emission. In the case of the red UV wing, there is a small allowed region (contour marked as EW=180 Å) for outward emission that corresponds to the covering factor of 38 %, but it appears only for low values of the ionizing photon flux. Thus, at least about object RM 102, we can conclude that the shielded sides of the clouds cannot be the dominant source of the Fe II emission in UV.
The strength of integrated Fe II emission in the optical band is overall higher when measured in terms of the EW and lower in terms of absolute flux (see Figures 4 and 5). Without microturbulence, only the low ionizing photon flux and relatively low-density lead to significant flux. Microturbulence again enhances the Fe II emission strength. The difference between the inward and outward emission is smaller than that for UV. Still, the non-irradiated sides of the clouds are not likely to emit enough radiation to explain the observed Fe II emission when the BLR covering factor of the order of 10 - 30% is considered. The small allowed region requires the covering of order of 50% for the total optical Fe II, it is only slightly broader for the red blend and never corresponds to the data from the object RM 102 for the blue blend.
Therefore, in further studies, we concentrate on the total Fe II flux and do not show separate contour plots for the outward radiation coming from the unilluminated cloud side.
3.2 Shape of UV Fe II emission
Baldwin et al. (2004) noticed two strong spikes at around 2600 and 2400 Å, resulting from UV 1 and UV 2+3 multiples, respectively, in the model-predicted spectra that are absent in the observed spectra (see Figure 6). In the standard photoionization model (including thermal velocity only), the optical depths at the illuminated surface of the cloud are small. Consequently, the energy deposited by the incident photons is radiated through the strongest resonance lines near 2400 and 2600 Å only that appear as the two spikes in the model spectrum of the cloud. The presence of these spikes in the model spectra is because Fe II mostly forms in the illuminated face of the cloud with significantly smaller optical depths, which permits the strongest resonance lines to exit directly. This issue of the difference in the shape of Fe II emission was largely investigated by Baldwin et al. (2004) and recently, by Sarkar et al. (2021). They suggested that including the microturbulence (100 km/s) into photoionization calculations can solve this shape problem. When microturbulence is incorporated into the model cloud, it increases Doppler broadening, which lowers the line optical depths. As a result, lines can escape more easily and appear brighter. Moreover, microturbulence increases the likelihood of continuum and wavelength coincident bound-bound pumping of levels upward (including the upper level of the original transition), since the line can now absorb photons over a larger Doppler width (Wills et al., 1985; Bottorff et al., 2000).
To examine this UV Fe II shape issue, we adopted the definition of the spike as the integrated Fe II flux over the wavelength ranges 2280–2430 Å and 2560–2660 Å and that of the gap as the integrated Fe II flux over 2430–2560 Å wavelength range from Baldwin et al. (2004) and computed spike/gap ratio for all the models. The contour plots for the spike/gap ratio for all the microturbulence velocities are shown in Figure 7. The observed value of the spike/gap ratio for the spectrum of RM 102 is 1.58 (see Table 2) and corresponds to the red dashed contours in Figure 7. When the microturbulence is zero, no such contour is seen. However, including microturbulence in our model calculations affects the shape of the UV Fe II and contours that match the measured value of the spike/gap ratio are visible, for a wide range of model parameters. So again, higher values of the microturbulence velocities are favoured.
In this study, we primarily examined six Fe II blends (listed in Table 1). The model parameter space in the nH- plane that predicts the EW (for a 100% covering factor) of Fe II consistent with the observed value differs for each blend. For instance, the parameter space for the optical Fe II broadband blend is more constrained than that for the UV Fe II broadband blend (see Figures 3 and 5). Therefore, in Figures 7–17, we have shaded only the region corresponding to the Fe II blend with the narrower parameter space among those Fe II blends used to calculate the ratio, except in Figures 7, 12 and 17. In Figure 7, the shaded region is consistent with the Fe II emission strength for the broadband UV Fe II blend, whereas in Figures 12 and 17 the shaded region specifically denotes the models that predict the observed optical blue Fe II for a covering factor of 100%.